OpenStudy (shicarak):

(2x 2 +5x−9)−(−4x+5)

OpenStudy (shicarak):

the 2 next to the 2x is supposed to be an exponent

OpenStudy (jalil.h):

@welshfella did i get it right

OpenStudy (welshfella):

what did you get?

OpenStudy (jalil.h):

196x +245x

OpenStudy (jalil.h):

but the 2 supposed to be small and at up top

OpenStudy (shicarak):

yes

OpenStudy (welshfella):

No @jalil.H

OpenStudy (welshfella):

do the second part as in last question:0 -(4x + 5) = -1(4x + 5) = ?

OpenStudy (welshfella):

sorry thats - (-4x + 5) = -1(-4x + 5)

OpenStudy (welshfella):

-1*-4x = what?

OpenStudy (jalil.h):

4x

OpenStudy (triciaal):

you can shift the number 6 to get the symbol for exponents

OpenStudy (welshfella):